OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8

OpenStudy (anonymous):

lets isolate x in the equation -x - y = -8 -x = y - 8 (multiply by -1 to make x positive) x = -y + 8 now sub -y + 8 in for x in the 1st equation

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

no thank you